OpenStudy (anonymous):

_____ 14. Tell whether the sequence is arithmetic, geometric, or neither. Find the next three terms of the sequence. Type your answer in the blank to the left. A. neither; 7, -20, 61 B. geometric;7, -20, 61 C. arithmetic; D. geometric;

hero (hero):

Given, r = {7, -20, 61...} Find a relationship between the first three terms of the sequence: The difference between 7 and -20 is \((-3)^3 = -27\) The difference between -20 and 61 is \((-3)^4 = 81\) Thus we can assume the next term of the sequence is \(61 + (-3)^5 = 61 - 243 = -182\) If \(-182\) is the next term of the sequence, then term following it is: \(-182 + ((-3)^6 = -182 + 729 = 549\) And the term following 549 is \(549 + (-3)^7 = 549 - 2187 = -1638\) In other words, the common difference between the terms is \((-3)^{n+2}\) Which is neither arithmetic nor geometric